The Earth may be considered to be a uniform sphere of radius 6.38 × 106 m. Its mass is assumed to be concentrated at its centre.


Given that the gravitational field strength at the Earth’s surface is 9.81 N kg–1, show that the mass of the Earth is 5.99 × 1024 kg.



1
Expert's answer
2022-02-07T08:42:51-0500

"F_1=F_2,"

"mg=\frac{GMm}{R^2},"

"g=\frac{GM}{R^2},"

"M=\frac{gR^2}G,"

"M=5.99\cdot 10^{24}~kg."